You are here: Foswiki>PartnerObserving Web>TechNightLogs>2009March20UT (21 Jul 2009, NormCushing)Edit Attach

20 March 2009 UT - LBC Technical PartnerObserving

Observers: O. Kuhn (@LBT), J. Hill (Tucson)
Telescope Operator: D. Gonzalez Huerta

Thin cirrus over all sky at sunset

~18:00 Both Red and Blue cameras filled by Doug. Red was last filled at 3pm, so 27 hours in between. It looks like it started to warm up at 17:30, so hold time is about 26.5 hrs

18:20 Turn on LBC - all looks OK.

18:27 Testing LBC preset for s/w group (new TCS build in first week of March, but not yet tested with LBC). Fast OB will slew to ra,dec = 7h,10d and take 1 sec exposure at each of 3 dither positions. David had to abort preset - waiting for OB to stop on its own, to avoid any problems from stopping an OB during slew (as during a preset to target before focus/collimation). My OB mistakenly observed only at blue.

18:34 OB is moving very slowly, but going along despite the fact that David had aborted preset. Nevertheless, it completed, and now we are moving on to try to take Tech chip sky flats.

18:40 Edited redchannel & bluechannel conf files to set MinimumScientificExposureTime = 0.5sec, so trackers will be activated during flats. Saved originals in ---20090320save Turning off and back on LBC.

18:40 Opening Dome.

18:47 Going to 1hr E, dec=25 to take flats. UY. No tech chip images are coming. Do I need to run lbckill/start? But I turned off/on LBC.

18:56 Turn off LBC. lbckill/start. lbcstart gives the following message:
[root@CMU lbccontrol]# ./lbcstart.sh
Starting LBC Control System: Binary Data Dictionary File /lbt/tcs/current/etc/./event.bin not found
Binary Data Dictionary File /lbt/tcs/current/etc/./event.bin not found
Server has been forked!
Looks like it worked, but why are these files not found?

BTurning back on LBC - why is LBC-Blue not coming up??? All systems "enabled" and not going to "initialization". Log says:
 
2009/03/20 02:01:07.231881 E B                   initialization not allowed on hinibited channel
shortly after channel inhibit was cleared
2009/03/20   02:01:08.123389           channel hinibit cleared

19:04 Turning off LBC again so as to get LBCBlue back. OK.

19:15 Taking gi flats for 10sec each. We are getting some lbcb tec chip images but not lbcr, because the minimum tec exposure is 16sec for red (> my 10sec exposure time) but only 8 sec for blue. There are stars on the flats, but I'll try 20sec flats next. Still not long enough for red, so increasing exp time to 40sec.

19:25 Clearing Active Optics & entering Z11=+1000 for SX. Encoder reset was done yesterday afternoon & MCSPU was not shut down since then.

19:29 Going to collimate. Both Blue & Red, necessary Z4 adjustment is too large, -12000 for both. But not sent, because of this. -12000 is below the threshold for printing any correction, all asterisks, but says it is sending data, and indeed SX and DX got -12000 for the initial correction.

(JMH: The *** problem has been fixed by changing the Z4 output format to I6.)

19:45 This first set of pupil images is very weird. Blue pupils have a 2nd hole at about 7 o'clock (is it the intra-focal 'squirrel'?) are are elongated along the lower left/upper right diag, and Red pupils are severely elongated right/left. David says no actuators are off - all are working fine.

19:51 DOFPIA suggestions:

A delay in data files appearing in Repository. Message says only
******* WARNING *******
Data files DID NOT appear in /Repository
Pls double check and then: press Q to exit or any other key to continue
But it would be helpful if it said which data files did not appear.

(JMH: DOFPIA doesn't know the names of the files it is waiting for, it is only waiting for two new images to appear. We extended the timeout from 63 to 99 seconds, so this message will be less likely to appear.)

Does DOFPIA stop after N iterations. It is going and going... so far 10 images.

(JMH: We added a command line parameter for DOFPIA to set the maximum number of iterations. MAXITER=xx)

20:00 Red is round, but Blue still elongated. Confirmed that there is no Z2/Z3 (i.e. mirror guiding) even though I still have minimum time for trackers set to 0.5sec, and guide images are being taken. Ask David to apply Z11=-334 to DX (-75 - 350*seeing) and Z11 = -720 (-300 -350*1.2) (seeing was 1.77 but I think it is fooled).

20:05 Try again with DOFPIA. Red nice, Blue is still very elongated. Red should have converged... but it has not sent the final Z11 which should have been -75-350*0.74=-334. [NOTE: this is because blue never converged. The new DOFPIA keeps iterating on Red, even though it has converged, so that it does not fall out of collimation while waiting for Blue (or vice versa)].

20:15 Clear AO on Blue and +1000 Z11 again. Give Z11 = -334 to Red.

20:18 Running DOFPIA again. Now Blue pupils are ROUND. Clearing AO did the trick. But it is not converging and the lower-left/upper-right elongation is returning. There is again the 2nd hole - like before it is along the spider, adjacent to the central hole.

20:36 Stopping DOFPIA, clearing SX AO & adding +1700 nm Z11. To DX give -334 Z11.

20:50 Running new dofpia. This version will allow user to
  • set threshold
  • max number iterations
  • red only or blue only
  • estimated Wavefront Image Quality (est WIQ) - value should be small compared to seeing
  • outputs also filter

21:00 Running dofpia again - both sides have values which are <400, but the convergence flag isn't raised.

21:50 Running the Z4/Z11 experiment (while John working on dofpia).
  • Run lbcfpia on blue/red until converged (battling my quick forgetting of the interactive nature of lbcfpia).
    • 045911, 045910
-227     85    200    -97    -74     52    n  1.02                   /Repository/lbcb.20090320.045911.fits diam= 63.1 pix
-330     94    137     21     57     66    n  0.56                   /Repository/lbcr.20090320.045910.fits diam= 63.8 pix
  • David adds Z11 = +1000nm to both SX and DX., 61.2 (B)
  • recollimate
-1987    147     83     -9      3   -634    y  0.91                   /Repository/lbcb.20090320.050309.fits diam = 61.2 (by eye on ds9)
  591   -124     84     30    -83   -173    y  1.20                   /Repository/lbcb.20090320.050735.fits
  154    378    -35     36    -61   -148    y  0.68                   /Repository/lbcb.20090320.051041.fits
  145   -173    -81    -35      5    -10    n  0.65                   /Repository/lbcb.20090320.051241.fits


-1858     83    157    175   -133   -374    y  0.43                   /Repository/lbcr.20090320.050308.fits diam  = 58.9 or 61.6 (by eye on ds9)
   74   -192     14   -144    173   -221    y  0.64                   /Repository/lbcr.20090320.050734.fits
   27     87   -117    -49    -63   -178    y  0.36                   /Repository/lbcr.20090320.051040.fits
  227     11     63     37      4    -40    n  0.39                   /Repository/lbcr.20090320.051241.fits

22:30 My measurements, +2 pix, indicate Z4 = -700nm (-350.3 nm Z4/pixel). We'll repeat experiment but this time subtract spherical: Z11=-1000nm.
  • Collimate blue and red again (using lbcfpia)
    
  0    128    122    -56     43     45    n  1.16  0.10                   /Repository/lbcb.20090320.054020.fits  V-BESSEL diam = 67.17
60    208     57    188     79    -43    n  0.65  0.14                   /Repository/lbcr.20090320.054019.fits  r-SLOAN    diam = 66.8
  • David giving Z11 = -1000 nm to both sides.
  • recollimate 054335(R),054336(B)
  1035     36    174    -98    -28    639    y  0.83  0.58                   /Repository/lbcb.20090320.054336.fits  V-BESSEL diam = 70pix
  -681     41   -295    133     19    271    y  0.84  0.34                   /Repository/lbcb.20090320.054657.fits  V-BESSEL
  -481   -217     10     40     14     59    y  0.92  0.22                   /Repository/lbcb.20090320.055046.fits  V-BESSEL
  -141    218    104    -15     30     39    y  0.87  0.12                   /Repository/lbcb.20090320.055318.fits  V-BESSEL


   849   -147    139      5   -101    522    y  0.55  0.48                   /Repository/lbcr.20090320.054335.fits  r-SLOAN diam = 69.5
  -624    203   -263     79     76    262    y  0.43  0.33                   /Repository/lbcr.20090320.054656.fits  r-SLOAN
  -389    188     88     54    -50     65    y  0.43  0.19                   /Repository/lbcr.20090320.055045.fits  r-SLOAN
  -219    -56     39      6     92     36    y  0.39  0.13                   /Repository/lbcr.20090320.055317.fits  r-SLOAN

23:00 Looks like -1000nm increased diam by 3 pixels. Running again, this time with increased Z11 gain in LBCFPIA.

  • Converged 061219,061220
    -7   -134   -163    -68    -40    -87    n  0.88  0.12                   /Repository/lbcb.20090320.061220.fits  V-BESSEL 64
    57     20   -162    -98     67    -33    n  0.44  0.11                   /Repository/lbcr.20090320.061219.fits  r-SLOAN 64.5
  • David sending Z11 = +1000 to both SX and DX
  • Recollimate: (note sign of Z11 and sign of Z4 identical - send Z11 = -1000nm, need to send Z4 ~ -1000nm to compensate).
 -1723   -223   -195     43     38   -667    y  0.85  0.81                   /Repository/lbcb.20090320.061617.fits  V-BESSEL 60
   811    277     52     45     11   -126    y  0.84  0.35                   /Repository/lbcb.20090320.061902.fits  V-BESSEL
   264   -113   -247     32   -118   -134    y  1.05  0.18                   /Repository/lbcb.20090320.062240.fits  V-BESSEL
    30    -34    302   -163    261   -203    y  1.06  0.22                   /Repository/lbcb.20090320.062635.fits  V-BESSEL


 -1810   -115    -57    194   -122   -305    y  0.39  0.75                   /Repository/lbcr.20090320.061616.fits  r-SLOAN  61   (about 3-3.5 pix smaller)
   592   -113    -24   -132    273   -354    y  0.43  0.36                   /Repository/lbcr.20090320.061901.fits  r-SLOAN
   510    187    219     79    -70   -162    y  0.54  0.26                   /Repository/lbcr.20090320.062237.fits  r-SLOAN
    83    251     20    227    149    -34    y  0.65  0.17                   /Repository/lbcr.20090320.062634.fits  r-SLOAN

  • Collimated at 062480

00:00 John edited dofpia to send Z4 with Z11.
  • It sends Z4 = 0.9*Z11.
  • Also you can set the threshold - 400nm is a reasonable value, default is 420.

00:00 testing it. But although all values < 420nm, program doesn't stop. Bug fixed - trying again.

00:16 Converged:
   540      0      0      0      0    600    y  0.00  0.00                     dofpia_sph3_adjustment
 -1551    345    353     71    225   -405    y  1.02  0.69                   /Repository/lbcb.20090320.070931.fits  V-BESSEL
   892    146   -271     31   -127    -48    y  0.95  0.38                   /Repository/lbcb.20090320.071126.fits  V-BESSEL
   307   -184    155   -148      7     25    y  1.09  0.17                   /Repository/lbcb.20090320.071321.fits  V-BESSEL
  -612      0      0      0      0   -681    y  0.00  0.00                     dofpia_sph3_adjustment

   540      0      0      0      0    600    y  0.00  0.00                     dofpia_sph3_adjustment
 -1185   -117     19   -113    121   -281    y  0.48  0.51                   /Repository/lbcr.20090320.070930.fits  r-SLOAN
   355    -42     73   -299    -43   -104    y  0.50  0.22                   /Repository/lbcr.20090320.071125.fits  r-SLOAN
   355    -49    -78    263    -72    -58    y  0.60  0.21                   /Repository/lbcr.20090320.071320.fits  r-SLOAN
  -257      0      0      0      0   -285    y  0.00  0.00                     dofpia_sph3_adjustment

00:16 RB_RV_superfoc (R and V Bessel - meant to use r-SLOAN as fastextra).
  V-Bessel Average best focus of 0.0163496 with FWHM of 3.41
  R-Bessel Average best focus of 0.117956 with FWHM of 2.78
00:25 dofpia exited with EOF error on input file, but file is correct size and displays OK. Data transfer may not have completed when idl wanted file.
Processing: /Repository/lbcr.20090320.072546.fits
FITS_READ ERROR: READU: End of file encountered. Unit: 101
         File: /Repository/lbcr.20090320.072546.fits

00:37 Converged - RB_rV_superfoc (r and V).
   Average best focus of 0.140948 with FWHM of 2.88 (vs 0.161 )
  Average best focus of 0.0149762 with FWHM of 3.33 (fit is not so good) 
00:46 Stopped OB while blue 'flushing'. Going to dofpia to check collimation. Run 'dofpia' but OB name has not changed - command rejected while another running. Had to cntl-c so David is giving Z11 = -600nm to both sides --- also -540nm Z4. Run dofpia again.

00:57 rV superfoc
 r Average best focus of 0.138403 with FWHM of 3.49
 V Average best focus of -0.0400098 with FWHM of 3.39
01:00 Ir superfoc
 r (LBCB)    Average best focus of 0.0439611 with FWHM of 3.08 (quite different from nominal value of 0.086)
 I (LBCR)   Average best focus of 0.0499653 with FWHM of 3.64
01:12 dofpia and then
  • RB_RUs_superfoc
R Average best focus of 0.113382 with FWHM of 3.70
SDT-U  Average best focus of 0.426446 with FWHM of 4.39 (not a good fit)
  • repeat RUs:
R  Average best focus of 0.12242 with FWHM of 3.13 (now OB has been centered better)
SDT-U Average best focus of 0.451741 with FWHM of 3.18 (good fit)
01:28 dofpia again then RB_ig_superfoc
i-SLOAN Average best focus of 0.103533 with FWHM of 2.27
g-SLOAN Average best focus of 0.082311 with FWHM of 2.92

z-SLOAN Average best focus of 0.0613662 with FWHM of 2.12
B-B Average best focus of 0.0630867 with FWHM of 2.69

I Average best focus of 0.0713499 with FWHM of 2.25
r Average best focus of 0.0800337 with FWHM of 2.64 (not a very good fit - smooth but not symm)
01:51 dofpia and then Ir followed by YU and F972-004mm_superfoc (w/ U-B for LBCB)
I  Average best focus of 0.0611233 with FWHM of 2.46
r  Average best focus of 0.0695461 with FWHM of 2.64

Y  Average best focus of 0.051126 with FWHM of 2.41
U Average best focus of 0.337409 with FWHM of 2.92

F972N20  Average best focus of 0.222545 with FWHM of 2.50 (exactly what is in the conf file).

02:11 dofpia then YU and zB
Y  Average best focus of 0.0585871 with FWHM of 2.29
U Average best focus of 0.333985 with FWHM of 2.65
 
z Average best focus of 0.0607938 with FWHM of 2.15
B Average best focus of 0.0704227 with FWHM of 2.67

F972 Average best focus of 0.203893 with FWHM of 2.28

02:35 new version of dofpia loaded - won't crash if no pupils found.

02:39 dofpia then ig
i Average best focus of 0.0895423 with FWHM of 2.15 (not a great fit)
g  Average best focus of 0.119168 with FWHM of 2.86 (again not a great fit)

02:55 We were at elev=44. Slewing back to higher elev, ~75deg

03:00 DOFPIA with RedLimit +dofpia,RedLimit=500+
i Average best focus of 0.0742534 with FWHM of 2.10
g --- not a good fit at all...

V (R)  Average best focus of 0.19249 with FWHM of 2.37
V (B) awful fit

R  Average best focus of 0.101707 with FWHM of 2.39
SU  Average best focus of 0.471885 with FWHM of 2.67 (good fit)
03:33 dofpia
i Average best focus of 0.0677246 with FWHM of 2.10
g Average best focus of 0.0655358 with FWHM of 3.34 (not a good fit - astig)

V (R) Average best focus of 0.189762 with FWHM of 2.44
V (B) ---
03:54 dofpia and then Ir and VV
I Average best focus of 0.0727754 with FWHM of 1.95
V (R)  Average best focus of 0.218363 with FWHM of 2.32
04:15 Slewing to a standard to take in-focus images at all Red filters - slew and stop immediately, then run dofpia.
  • first with old focus offsets
  • fwhm: r, R, I, Y, V, i, z, F = 1.91-1.98, 2.4, 1.9, 1.9, 2.5 2.2-2.5, 2.4, 1.56 (2.2 r on 2nd dither)
  • aborted OB, turn off/on camera to pick up new filter focus offsets
  • dofpia again
  • a collimation point at 55deg elev: 115128 both blue/red
  • fwhm r, R, I, Y, V, i, z, F = 2.3, 2.3, 2.1, 1.9-2.3(coma?), 2.6, 2.35, 2-2.3, 1.6-2, (2.4 r on 2nd dither)
    • this 2nd set is worse, but it is hard to distinguish whether it is bad collimation - all image shapes look off - or focus.

05:05 Collimation points (T_smt = 4.6C) SX and DX GOs are all zero except for: Z = -1mm
  • 55 deg 115128, 115128
  • 83 deg 121101, 121058
  • 75 deg 122001, 122001
  • 65 deg 122643,122642
  • 54 deg 123243, 123242
  • 44 deg getting too light for fpia to find pupils (123712). Cntl-C.

05:43 Closing up. Looks like there is still cirrus over all sky at twilight.

-- OlgaKuhn - 20 Mar 2009
Topic revision: r6 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback